# A Computational Theory and Semi-Supervised Algorithm for Clustering

### v1

@@ Abstract
A computational theory for clustering and a semi-supervised clustering algorithm is presented. Clustering is defined to be the obtainment of groupings of data such that each group contains no anomalies with respect to a chosen grouping principle and measure; all other examples are considered to be fringe points, isolated anomalies, anomalous clusters or unknown clusters. More precisely, after appropriate modelling under the assumption of uniform random distribution, any example whose expectation of occurrence is $<1$ with respect to a group is considered an anomaly; otherwise it is assigned a membership of that group. Thus, clustering is conceived as the dual of anomaly detection. The representation of data is taken to be the Euclidean distance of a point to a cluster median. This is due to the robustness properties of the median to outliers, its approximate location of centrality and so that decision boundaries are general purpose. The kernel of the clustering method is Mohammad’s anomaly detection algorithm, resulting in a parameter-free, fast, and efficient clustering algorithm. Acknowledging that clustering is an interactive and iterative process, the algorithm relies on a small fraction of known relationships between examples. These relationships serve as seeds to define the user’s objectives and guide the clustering process. The algorithm then expands the clusters accordingly, leaving the remaining examples for exploration and subsequent iterations. Results are presented on synthetic and realworld data sets, demonstrating the advantages over the most widely used clustering methods.
@@ 1 Introduction
In the present work both the unsupervised and semi-supervised approaches to clustering are of interest—leaving aside a strict definition. The former has benefits of ease of use while the latter can use explicit prior information from the user to guide or constrain the clustering to solutions that are desired amongst the vast number of possibilities.
@@ 1.1 Prior Art
For example, in hierarchical clustering such as the agglomerative method [11], users are required to specify the granularity at which clusters are to be taken. This has the advantage that clustering can be considered at different resolutions or levels but requires nontrivial (and perhaps unintuitive) specification on the users part. There are also a number of linkage methods to select from that can give different results. The runtime of the algorithms and storage requirements are also prohibitively expensive with increasing number of examples; being $O(n^{2}logn)$ and $O(n^{2})$, respectively. Another important aspect to note is that all merges are final in the building of clusters. This can be problematic when for example there are anomalies present since the grouping cannot adjust or correct later to a better solution. Furthermore, the hierarchical method does not natively associate a score with an observation or predict the cluster label of a new observation given the current clustering, this restricts its utility somewhat unless additional classifiers are applied on top.
@@
Density based methods work on the premise that clustering locates regions of high density that are separated from one another by regions of low density. The most popular method is DBSCAN [3] which can cluster arbitrarily shaped data distributions efficiently ($O(nlogn)$ runtime and in $O(n)$ memory), while also potentially detecting anomalies. However, it requires the specification of parameters that are highly sensitive and difficult to set in practice even with the ‘elbow’ identification method. Thus, it can fail to give good results in even what appear to be simple data sets if clusters are too close to one another or there are large differences in densities. The method also benefits from post-processing to remove small groups of data that may get incorrectly classed as clusters, and it has been found in private experiments that its anomaly detection results are unsatisfactory. DBSCAN also does not natively provide a cluster association score for every observation or have predictive ability over new observations. The only way to handle the latter issue is to rerun the clustering again with the new data.
Probability based models fit statistical models to the data and estimate parameters, with the mixture of gaussians method being the most popular. In the cases where the assumptions are met it can give very good results, and the data points are not only assigned a grouping but each also has an associated probability score of belonging. The use of distributions enables simple characterisation of the clusters produced by the parameters, and has some general applicability in terms of handling ellipsoidal clusters. However, the method can be particularly slow due to the use of the Expectation Maximisation algorithm for finding the distribution parameters and hence is not practical for large data sets. It also does not work well when there is little data, if the data points are nearly co-linear, when the number of groups is assumed incorrectly, or when the gaussian data distribution assumption is not met. Noise and anomalies are also not natively handled since it computes a partitioning.
@@
Of all the clustering algorithms mentioned thus far, it is still the simple method of $k$-means [5, 4] that is the most widely known and stands out as the algorithm of choice for practitioners—despite it being over $50$ years old. It does however have several drawbacks which has resulted in much additional development, but most of which have not gained traction for practical use because of the additional complexities they introduce without necessarily better results. Due to it being a centroid based method, and because of its widespread appeal and use, comparisons in the present paper will focus more on $k$-means since any algorithm that can better or complement it will be a significant contribution to the field of cluster analysis. Hence, this algorithm will be elaborated upon.
@@
However, $k$-means is sensitive to the initialisation centroids, resulting in getting stuck in local optima so that multiple runs with random initialisation is sometimes recommended. However, this may not necessarily work well depending on the data and number of clusters specified. Thus, another approach includes taking samples of points and clustering them using hierarchical clustering where the centroids of those clusters are used as seeds. This approach may only work well if the sample size is small due to the computational complexity of hierarchical clustering and if $k$ is relatively small compared to the sample size. Note that the number of clusters still requires specification and its empirical success remains unclear. Other extensions include $k$-means+++++ + and bisecting $k$-means. The former picks centroids incrementally where new centroids are taken to be a random one that is far from its closest centroid and based on a probability measure that is proportional to the square of its distance. The latter splits the data into two clusters, and repeatedly operates on one of the clusters until $k$ clusters have been produced. Interestingly this method also yields a hierarchical clustering by recording the sequence of clusterings produced as $k$-means bisects each cluster.
@@
The optimisation criteria and the computation of centroids using local means makes $k$-means sensitive to anomalies which can lead to suboptimal partitions of the data. However, detecting and removing anomalies first is a challenging problem in of itself. In the case of clustering followed by anomaly removal, the sequential combination requires further analysis and practical investigation due to the distorting affect anomalies can have on the initial clustering. Another couple of methods similar to $k$-means, called $k$-medoids and $k$-medians have been proposed to overcome the problem of sensitivity to anomalies. The former takes an actual point to represent a cluster rather than the mean, while the latter uses medians as representatives. However, both are generally computationally slower than $k$-means, particularly when the data set and $k$ are very large.
@@
Closely related to the problem of anomalies is that $k$-means can give undesirable partitions in the presence of clusters with different sizes or densities. Additionally, because it is a partitioning algorithm it will always place an example into a cluster even if it is an anomaly, a cluster of anomalies or in fact belongs to another unknown cluster. Thus, it is important to differentiate between algorithms that return clusters (such as DBSCAN) and partitioning algorithms since the term clustering is overloaded. It is usually the case that when users want to carry out cluster analysis, they desire groups of data that are similar to be returned while excluding that which does not belong to the group. This has implications with regard to what should be computed in the problem of clustering data and whether $k$-means is merely approximating what is desired of an algorithm.
@@
$k$-means has also been the basis of many semi-supervised clustering algorithms, with perhaps the most well known of them being $COPK$-means [10] and $PCK$-means [1]. The former operates by assimilating hard constraints from the user that cannot be violated while carrying out $k$-means clustering. Thus, the constraint labelling is assumed noiseless. Similarly, the latter method incorporates constraints from the user, but these are soft constraints where the algorithm suffers a penalty for violation. Both algorithms have claimed to improve clustering results over the standard unsupervised version. However, although the application domain is important (since oftentimes some labelled data or constraints are available), these methods have received little attention in practice. This is perhaps because the data sets used in the original works are relatively small, the algorithms slower to run, and because the results are not substantially improved. Furthermore—if it is of concern—most of the practical drawbacks of the $k$-means algorithm are also carried over into these semi-supervised implementations.
@@ 1.2 Contributions
The present work builds a clustering algorithm from the ground up by following the above analyses and using human visual perception as inspiration. This approach was successfully used by Mohammad [8] in his development of anomaly detection, which will form a core component of the clustering method. Indeed, clustering is considered the complement of anomaly detection (sensing distinctness), where clusters are deemed to be sets of data points that do not contain any anomalies with respect to a chosen grouping principle and measure. Mohammad’s definition of an anomaly is used, and data points are represented as the distance to the median of the parent group so that hyperspherical clusters are formed in order to have a solution that is generally applicable and practically useful like the $k$-means algorithm (while accepting failure in other and non-hyperspherical distributions). Furthermore, anomaly detection being the core of the approach naturally results in isolated points or clusters being labelled as anomalies so that clustering—and not a partitioning of the data—is carried out.
@@
The clustering method presented in this paper is parameter-free but assumes an interactive process between human and machine, where users must provide a small amount of labelled data that approaches a constant with increasing size of the data per cluster, so that seed members of known groups are obtained. This little information is enough to constrain and guide the algorithm to grow meaningful clusters and carry out both local and global anomaly detection. As users interact with the clustering outputs, further observations can be sampled, analysed and discovered, while existing clusters may be modified. The method—being based on Mohammad’s anomaly detection algorithm—is fast, intuitive and light on memory usage. Furthermore, because there is no best clustering algorithm in general, a framework is also provided (but not yet developed) for dealing with other useful representations of data in order to tackle different data distributions.
To detail the contributions of this work, the rest of the paper is laid out as follows: Section 2 provides a computational theory of clustering and describes the parameter-free, semi-supervised clustering algorithm. Section 3 gives some empirical results on synthetic one-dimensional and two-dimensional data sets to intuitively understand the performance of the algorithm. Additional results on publicly available data sets are also provided to demonstrate realworld applications. Finally, section 4 concludes the paper and provides some thoughts for future work.
@@ 2.1 Tri-level Hypothesis
The third level of analysis, known as the implementation level, focuses on the physical realisation of the computations. This can involve various systems, ranging from the human brain to specialized mechanical machines, but for our purposes it will be a single computer. Nevertheless, it may be important to consider the potential extensions to distributed implementations in scenarios such as big data applications and distributed devices.

### v2

@@ Abstract
A computational theory for clustering and a semi-supervised clustering algorithm is presented. Clustering is defined to be the obtainment of groupings of data such that each group contains no anomalies with respect to a chosen grouping principle and measure; all other examples are considered to be fringe points, isolated anomalies, anomalous clusters or unknown clusters. More precisely, after appropriate modelling under the assumption of uniform random distribution, any example whose expectation of occurrence is $<1$ with respect to a group is considered an anomaly; otherwise it is assigned a membership of that group. Thus, clustering is conceived as the dual of anomaly detection. The representation of data is taken to be the Euclidean distance of a point to a cluster median. This is due to the robustness properties of the median to outliers, its approximate location of centrality and so that decision boundaries are general purpose. The kernel of the clustering method is the perception anomaly detection algorithm, resulting in a parameter-free, fast, and efficient clustering algorithm. Acknowledging that clustering is an interactive and iterative process, the algorithm relies on a small fraction of known relationships between examples. These relationships serve as seeds to define the user’s objectives and guide the clustering process. The method then expands the clusters accordingly, leaving the remaining examples for exploration and subsequent iterations. Results are presented on synthetic and realworld data sets, demonstrating the advantages over the most popular unsupervised and semi-supervised clustering methods.
@@ 1 Introduction
In the present work both the unsupervised and semi-supervised approaches to clustering are of interest—leaving aside a strict definition. The former usually has benefits of ease of use while the latter can use explicit prior information from the user to guide or constrain the clustering to solutions that are desired amongst the vast number of possibilities.
@@ 1.1 Prior Art
For example, in hierarchical clustering such as the agglomerative method [29], users are required to specify the granularity at which clusters are to be taken. This has the advantage that clustering can be considered at different resolutions or levels but requires nontrivial (and perhaps unintuitive) specification on the users part. There are also a number of linkage methods to select from that can give different results. The algorithms’ runtime and storage requirements are often prohibitively expensive in practice, with time complexity scaling as $O(n^{2}logn)$ and space complexity as $O(n^{2})$, as the number of examples increases. This is largely due to the need to store and repeatedly update the full distance matrix after each cluster merge, which adds significant computational overhead. Another important aspect to note is that all merges are final in the building of clusters. This can be problematic when for example, there are anomalies present, since the grouping cannot adjust or correct later to a better solution. Furthermore, the hierarchical method does not natively associate a score with an observation or predict the cluster label of a new observation given the current clustering, this restricts its utility somewhat unless additional classifiers are applied on top.
@@
Density based methods work on the premise that clustering locates regions of high density that are separated from one another by regions of low density. The most popular method is DBSCAN [9], which can efficiently cluster arbitrarily shaped data distributions with $O(nlogn)$ runtime and $O(n)$ memory, in typical low-to-moderate dimensional settings, while also potentially detecting anomalies. However, in high-dimensional data, DBSCAN can become significantly slower—often approaching $O(n^{2})$ complexity-due to the increased computational cost of neighborhood queries, as finding $\varepsilon$-neighbors becomes less efficient and more computationally expensive. Moreover, it requires the specification of parameters that are highly sensitive and difficult to tune in practice even when using methods such as the ‘elbow’ method for parameter selection. Consequently, it can fail to give good results in even what appear to be simple data sets if clusters are too close to one another or there are large differences in densities. The method also benefits from post-processing to remove small groups of data that may get incorrectly classed as clusters, and it has been found in private experiments that its anomaly detection results can be unsatisfactory (see Figures 5 and 11 for examples). Additionally, DBSCAN does not natively provide a cluster association score for every observation nor does it offer predictive ability over new data points; the latter being handled by rerunning the algorithm again with new data.
Probability-based models aim to fit statistical distributions to the data and estimate their parameters, with Gaussian Mixture Models (GMMs) being among the most widely used. When the underlying assumptions are met, GMMs can yield strong clustering performance, providing not only group assignments but also probabilistic scores indicating the degree of association between data points and clusters. By modelling clusters as distributions, GMMs allow for interpretable, ellipsoidal cluster structures and offer a degree of flexibility across different types of data. However, the method relies on the iterative Expectation-Maximisation (EM) algorithm to estimate the distribution parameters, which can be computationally demanding. Although the cost of each EM iteration is approximately $O(nkd^{2})$—with $n$ the number of data points, $k$ the number of mixture components, and $d$ the dimensionality—the total number of iterations needed for convergence is highly variable and can grow with data complexity. As a result, the overall runtime is often substantial, particularly for large or high-dimensional datasets. The quadratic scaling with dimension further exacerbates this issue. In addition, GMMs are sensitive to the choice of $k$, and their performance can degrade when data is poorly separated, colinear, limited in quantity, or deviates from Gaussian assumptions. Noise and anomalies are also not natively handled, as GMMs produce a hard partitioning of the data.
@@
Of all the clustering algorithms mentioned thus far, it is still the simple method of $k$-means [19, 10] that is the most widely known and stands out as the algorithm of choice for practitioners—despite it being over $50$ years old. It does however have several drawbacks which has resulted in many extensions, but most of which have not gained traction for practical use because of the additional complexities they introduce without necessarily significantly better results. Since it is a centroid based method, and because of its widespread appeal and use, comparisons in the present paper will lean towards $k$-means because any algorithm that can better or complement it will be a significant contribution to the field of cluster analysis.
@@
However, $k$-means is sensitive to the initialisation centroids, resulting in it getting stuck in local optima so that multiple runs with random initialisation is sometimes recommended. However, this may not necessarily work well depending on the data and number of clusters specified. Thus, another approach includes taking samples of points and clustering them using hierarchical clustering where the centroids of those clusters are used as seeds. This approach may only work well if the sample size is small due to the computational complexity of hierarchical clustering and if $k$ is relatively small compared to the sample size. Note that the number of clusters still requires specification and its empirical success remains unclear. Other important and useful extensions include $k$-means+++++ + and bisecting $k$-means. The former picks centroids incrementally where new centroids are taken to be a random one that is far from its closest centroid and based on a probability measure that is proportional to the square of its distance. The latter splits the data into two clusters, and repeatedly operates on one of the clusters until $k$ clusters have been produced. Interestingly this method also yields a hierarchical clustering by recording the sequence of clusterings produced as $k$-means bisects each cluster.
@@
The optimisation criteria and the computation of centroids using local means makes $k$-means sensitive to anomalies which can lead to suboptimal partitions of the data. However, detecting and removing anomalies first is a challenging problem in of itself. In the case of clustering followed by anomaly removal, the sequential combination requires further analysis and practical investigation due to the distorting effect anomalies can have on the initial clustering. Another couple of methods similar to $k$-means, called $k$-medoids and $k$-medians have been proposed to overcome the problem of sensitivity to anomalies. The former takes an actual point to represent a cluster rather than the mean, while the latter uses medians as representatives. However, both are generally computationally more expensive than $k$-means without significant improvement in results.
@@
Closely related to the problem of anomalies is that $k$-means can give undesirable partitions in the presence of clusters with different sizes or densities. Additionally, because it is a partitioning algorithm it will always place an example into a cluster even if it is an anomaly, a cluster of anomalies or in fact belongs to another unknown cluster. Thus, it is important to differentiate between algorithms that return clusters (such as DBSCAN) and partitioning algorithms since the term clustering is overloaded. It is usually the case that when users want to carry out cluster analysis, they desire groups of data that are similar to be returned while excluding that which does not belong to the group. This has implications with regard to what should be computed in the problem of clustering data and whether $k$-means is merely approximating what is desired of an algorithm. The partioning problem is also apparent when the dimensionality of a data set is $>=10$ and the number of desired clusters is $k>=20$; in this situation the $k$-means algorithm often converges with one or more clusters which are either empty or contain very few data points [6].
@@
$k$-means has also been the basis of many semi-supervised clustering algorithms, with perhaps the most popular being constrained $k$-means [6], seeded $k$-means [2], $COPK$-means [28], $PCK$-means [3] and MPCKMeans [5]. These algorithms work in a number of different ways. Constrained $k$-means explicitly adds $k$ contraints to the underlying clustering optimisation problem requiring that each cluster have at least a minimum number of points in it. Seeded $k$-means initialises the centroids using pre-defined seeds, which can improve the clustering results when domain knowledge is available. A variant of this we call constrained seeded $k$-means fixes the supplied seeds and runs clustering on the non-seed data. $COPK$-means operates by assimilating hard constraints from the user (e.g., must-link and cannot-link) that cannot be violated while carrying out $k$-means clustering. Thus, the constraint labelling is assumed noiseless. $PCK$-means also incorporates constraints from the user, but these are soft constraints where the algorithm suffers a penalty for violation. Finally, MPCKMeans enhances the constrained clustering approach by combining soft pairwise constraints with metric learning so that individual metrics (Mahalanobis distance) for each cluster are learned which permits clusters of different shapes.
Other semi-supervised clustering algorithms of interest include semi-supervised DBSCAN, which uses labelled data to help obtain suitable parameters that characterise $each$ of the different clusters [16]. Semi-supervised Gaussian Mixture Models (GMMs) adapt the probabilistic clustering framework by integrating label information—either in the form of partially labelled data or pairwise constraints [26]. Semi-supervised spectral clustering enhances traditional spectral clustering by integrating prior knowledge, such as pairwise constraints or labelled data, into the similarity graph or its Laplacian matrix. This is commonly done by modifying the affinity matrix or adapting the graph partitioning objective to satisfy the imposed constraints [32, 13].
All the semi-supervised algorithms discussed thus far have claimed to improve clustering performance over the underlying unsupervised base versions. However, although the application domain is important (since oftentimes some labelled data or constraints are available), these methods have received less attention in practice. This is likely due to several factors: many of the original studies were conducted on relatively small datasets with a lack of extensive benchmarking against strong baselines; the algorithms themselves tend to be computationally expensive, especially on medium-to-large datasets; and available implementations are often research-grade, with few mature, well-supported libraries for practical use. Moreover, the performance gains are frequently modest unless an impractically large number of constraints is provided. Finally, any practical limitations inherent in the base clustering algorithm typically persist in its semi-supervised variants, further limiting their appeal in real-world applications.
In recent years, deep learning-based clustering methods have garnered significant interest for their ability to handle complex, high-dimensional, and unstructured data. Unlike traditional clustering techniques, which typically operate on the original feature space or rely on manually engineered features, deep learning models can learn more meaningful data representations prior to clustering [27, 24]. Notably, substantial performance gains are often achieved when feature extraction and cluster assignment are jointly optimised within an end-to-end framework. A seminal example of this approach is Deep Embedded Clustering (DEC) [30], which first uses a deep autoencoder to learn a low-dimensional latent representation of the data. The decoder is then discarded, and the encoder is fine-tuned by simultaneously refining the latent features and optimising cluster assignments. This is done by minimising a clustering-oriented loss, typically based on the Kullback-Leibler (KL) divergence between a soft cluster assignment distribution and a target distribution that encourages more confident cluster assignments. Another method, Deep Clustering Network (DCN) [31], jointly optimises a deep autoencoder and a k-means clustering objective within the latent space, balancing reconstruction accuracy and clustering performance. More recently, SHADE (Structure-preserving High-dimensional Analysis with Density-based Exploration) [4] has extended deep clustering by introducing a density-connectivity loss that enables the discovery of arbitrarily shaped, non-Gaussian clusters. It automatically detects both clusters and noise without needing predefined cluster counts, making it appealing in complex, unsupervised scenarios. While deep clustering methods have the potential to outperform traditional clustering approaches, this advantage must be validated through comprehensive experimental evaluation, since they often still utilise unsupervised learning algorithms. Moreover, any performance gains are often offset by increased model complexity, higher computational cost, and a lack of robust, production-ready implementations.
Semi-supervised adaptations of deep learning-based clustering methods have also been proposed, typically incorporating constraints into the objective function to leverage prior knowledge and guide the clustering process [12]. For instance, Semi-Supervised Deep Embedded Clustering (SSDEC) extends DEC by integrating pairwise constraints [25]. Another method, Deep Constrained Clustering (DCC) [33], provides a flexible framework that can incorporate various types of constraints — including pairwise relations, instance-level difficulty, and cluster size balance — by employing differentiable surrogate loss functions. In contrast, SHADE-CC [11] is not based on deep learning; instead, it leverages evolutionary optimisation, specifically the Success-History based Adaptive Differential Evolution algorithm, to identify high-quality clusterings under constraints. While evolutionary methods like SHADE-CC are generally not designed for large-scale data, they can be effective on small to medium-sized datasets with complex constraint structures.
Finally, it is worth noting that several works have attempted to combine clustering with outlier detection, although these tasks are often treated independently. This separation can be problematic, as outliers can substantially distort cluster structures, while reliably identifying outliers remains a challenging problem in itself. To address this, some methods seek to integrate outlier detection directly within the clustering process. A well-known example is DBSCAN, which naturally identifies low-density points as noise or outliers while forming clusters. However, in DBSCAN, outlier detection is a secondary outcome rather than a primary objective, and its effectiveness can be inconsistent depending on parameter settings and data characteristics. An alternative approach is k-means–(minus minus) [7], an extension of the standard k-means algorithm that iteratively excludes a fixed number of points considered outliers based on their distance to the nearest cluster centroid. While this improves robustness, it introduces an additional parameter—the number of outliers to remove per iteration—which (as stated by the authors) is very difficult to determine in practice. Another notable method, Clustering with Outlier Removal (COR) [18], jointly optimises clustering and outlier detection by performing multiple runs of k-means–. These runs produce varied partitions, which are then used to map data points into a binary space, improving resilience to outliers. The final clustering is obtained by optimising a unified objective based on Holoentropy. Although COR has claimed good performance, it also requires careful parameter tuning and can be computationally expensive on large datasets due to the repeated k-means– runs, binary transformations, and optimisation procedure.
@@ 1.2 Contributions
The present work builds a clustering algorithm from the ground up by following the above analyses and using human visual perception as inspiration. This approach was successfully used by Mohammad [22] in his development of the anomaly detection algorithm called perception, which will form a core component of the clustering method. Indeed, clustering is considered the complement of anomaly detection (sensing distinctness), where clusters are deemed to be sets of data points that do not contain any anomalies with respect to a chosen grouping principle and measure. Mohammad’s definition of an anomaly is used, and data points are represented as the distance to the median of the parent group so that hyperspherical clusters are formed in order to have a solution that is generally applicable and practically useful like the $k$-means algorithm (while accepting failure in other and non-hyperspherical distributions). Furthermore, anomaly detection being the core of the approach naturally results in fringe points, isolated points or small groups being labelled as anomalies so that clustering—and not a partitioning of the data—is carried out.
@@
The clustering method presented in this paper is parameter-free but assumes an interactive process between human and machine, where users must provide a small amount of labelled data that approaches a constant with increasing size of the data per cluster, so that seed members of known groups are obtained. This little information is enough to constrain and guide the algorithm to grow meaningful clusters and carry out both local and global anomaly detection. As users interact with the clustering outputs, further observations can be sampled, analysed and discovered, while existing clusters may be modified. The method—being based on the perception algorithm—is fast, intuitive and light on memory usage. Furthermore, because there is no universally best clustering algorithm, a framework is sketched for dealing with other useful representations of data in order to tackle different data distributions.
To detail the contributions of this work, the rest of the paper is laid out as follows: Section 2 introduces the computational theory underpinning the approach and presents the proposed parameter-free, semi-supervised clustering algorithm. Section 3 illustrates the algorithm’s performance on synthetic one- and two-dimensional datasets to provide an intuitive understanding of its behaviour. Additional results on publicly available data sets are also provided to demonstrate realworld applications. (Readers are encouraged to visit the GitHub repository linked in the footnote on the first page of the present paper to explore additional examples and experiment with the method.) Finally, Section 4 concludes the paper and outlines directions for future work.
@@ 2.1 Tri-level Hypothesis
The third level of analysis, known as the implementation level, focuses on the physical realisation of the computations. This can involve various systems such as the human brain or specialized mechanical machines; but for our purposes it will be a single computer. Nevertheless, it may be important to consider the potential extensions to distributed implementations in scenarios such as big data applications and distributed devices.